Suppose that the true distribution g(y) generating data and the specified model f (y) have normal distribution N(m, t2) and N(u, o²), respectively.


(a) Show that


Ec[logg (Y)] = -log(2rt²) - 1


1


where EG[] is an expectation with respect to the true distribution N(m, t²).


1
Expert's answer
2021-05-07T14:18:51-0400

The Normal distribution of generating data g(y) and specified model f(y) is-

g(y)~N("m,t^2")

f(y)~N"(\mu,o^2)"



The probability density function is-


"Y=\dfrac{1}{\sqrt{2\pi}}e^{\frac{-x^2}{2}}"


"logY=log\dfrac{1}{\sqrt{2\pi}}+loge^{\frac{-x^2}{2}}=log\dfrac{1}{\sqrt{2\pi}}+\dfrac{-x^2}{2}~~~~~-(1)"



So "E(logY)=\int_{0}^{r}xlogYdx"


     "=\int_{0}^{r} xlog\dfrac{1}{\sqrt{2\pi}}+\dfrac{-x^3}{2}dx\\[9pt] =\dfrac{x^2}{2}log\dfrac{1}{\sqrt{2\pi}}|_0^r-\dfrac{x^4}{8}|_0^r\\[9pt]=\dfrac{r^2}{2}log\dfrac{1}{\sqrt{2\pi}}-\dfrac{r^4}{8}\\[9pt]=-log(2rt^2)-1"


Hence "E(logY)= -log(2rt^2)-1"